OpenStudy (anonymous):

tell whether the sequence 1/3, 0, 1, -2.... is arithmetic, geometric, or neither. find the next three terms o the sequence.

OpenStudy (anonymous):

u sure

OpenStudy (amistre64):

does it follow any properties for arith or geo?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

how can we chk if it does or not tho?

OpenStudy (anonymous):

I AM NOT SURE

OpenStudy (amistre64):

an arith seq has the same number added to each term to get the next one. your first term is: 1/3 your second term is: 0 1/3 + d = 0 what would d have to be in this case?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

1/3 + 0 does not equal 0

OpenStudy (anonymous):

OK SORRY

OpenStudy (amistre64):

1/3 + d = 0 ; solve for d by subtracting 1/3 from each side -1/3 -1/3 ------------- d = -1/3

OpenStudy (amistre64):

now that we should have a proper "d", lets see if it fits 0 + d = 1 0 - 1/3 = 1 is false; so this cant be an arith seq can it

OpenStudy (anonymous):

SO ITS NOT GOING TO BE THAT THE U HAVE TO SEE IF ANYTHIG MULTIPLYS FOR GEOMETRIC

OpenStudy (amistre64):

correct 1/3 * d = 0 ; only when d=0 0 * r = 1 , well, never .... so it doesnt hold to a geo seq either

OpenStudy (anonymous):

SO ITS NOT GOING TO BE THAT I ALREADY TRIED THAT SO THE ANSWER IS NEITHER

OpenStudy (anonymous):

THANKS SOMUCH FOR SHOWING ME THAT

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

THANKS